OpenStudy (anonymous):

Rewrite the following expressions in x=my + b form. a. 3x + 6y + 9 = 0 b. -2x + 6y - 4 = 0 c. -4x - 2y = 0

jhonyy9 (jhonyy9):

a. x=(-6y-9)/3=-2y-3 b. x=(6y-4)/2=3y-2 c. x= 2y/-4 = -y/2